OpenStudy (anonymous):

50g of steel wire was dissolved in excessive of diluted sulphuric acid and the solution was made up to 250ml in a standard flask. 25ml of this solution was pipetted into a conical flask and needed 25.45ml of 0.02M KMnO4 for complete oxidation. calculate the percentage of iron in the sample.

OpenStudy (anonymous):

apa??

OpenStudy (anonymous):

Your first equation: Fe+H2SO4->FeSO4+ H2 When you add KMnO4, n factor for KMnO4=5(Mn7+ to Mn2+) and for Fe it is 1(Fe2+ to Fe3+) so by law of equivalents you have 5*0.02*25.45=x*1*25. (x is molarity of FeSO4). x=2.545/25=0.1018 No of moles in 250ml or 1/4 L= 0.1018*1/4=0.02545. Moles of FeSO4=moles of Fe. Moles of Fe in the wire=0.02545. Weight=moles*molecular weight=0.02545*55=1.4g. Now find precentage of iron. (I have a feeling i made a mistake in calculations- i always do. The method is correct).